OpenStudy (anonymous):

x/2x+2-2x/4x+4+2x-3/x+1 solve the equation

OpenStudy (anonymous):

\[\frac{ x }{2x+2 }-\frac{ -2x }{ 4x+4 }+\frac{ 2x-3 }{ x+1 }\]

OpenStudy (gorv):

4x+4 in this term can we take 2 common

OpenStudy (gorv):

??

OpenStudy (anonymous):

4(x+1)

OpenStudy (gorv):

yes but just take only 2 out not 4

OpenStudy (anonymous):

2(2x+2)

OpenStudy (gorv):

yeah righy

OpenStudy (gorv):

\[\frac{ x }{ 2x+2 }-\frac{ -2x }{ 2*(2x+2) }+\frac{ 2x-3 }{ x+1}\]

OpenStudy (gorv):

in second term can we cancel 2 ??

OpenStudy (anonymous):

I don't get when you said we can cancel 2 . from this I must get a LCD to get rid of my fractions right

OpenStudy (gorv):

\[\frac{ -2x }{ 2(2x+2)}\]

OpenStudy (gorv):

in this term 2 in denominator will cancel or cut or divide the 2 in numerator agree??

OpenStudy (anonymous):

yes true I agree which would leave neg x over 2x+2

OpenStudy (gorv):

there 2 continuous negative sign so it will become positive agree??

OpenStudy (anonymous):

yes

OpenStudy (gorv):

\[\frac{ x }{ 2x+2 }+\frac{ x }{ 2x+2}=??\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so my lcd is 2x+2 and x+1

OpenStudy (gorv):

\[\frac{ 2x }{ 2x+2 }+\frac{ 2x-3 }{ x+1 }\]

OpenStudy (gorv):

take 2 common from 2x+2

OpenStudy (anonymous):

2(x+1)

OpenStudy (gorv):

yeah right now can we cancel 2 ??

OpenStudy (anonymous):

true I noticed that as before with th e secon term right

OpenStudy (gorv):

yeah

OpenStudy (gorv):

so after canceling 2

OpenStudy (gorv):

\[\frac{ x }{ x+1 }+\frac{ 2x-3 }{ x+1 }\]

OpenStudy (gorv):

hope you can solve it now

OpenStudy (anonymous):

wheres the third term ?

OpenStudy (gorv):

@sam87211

OpenStudy (anonymous):

if you sum it up I learn better by seeing the steps im a bit confused because its spced out